# Multifractal critical phase driven by coupling quasiperiodic systems to electromagnetic cavities

### v1

@@ III Results
We begin our analysis by discussing the established properties of the AAH model without a cavity, which will serve as a baseline when discussing the effects of the coupling with photons. Figure 2 (a) shows the spectrum of the model for different disorder strengths. Notice that all eigenstates change their IPR near $g_{AA}=2$, as demanded by the self-dual property of the model. To probe the critical region it is more convenient to examine the $\overline{\text{IPR}}$ for different system sizes, as shown in Fig. 3 (a). Notice that for $g_{AA}>2$ $\overline{\text{IPR}}$ is large and weakly dependent on the system size, whereas for $g_{AA}<2$ it is small and decreases as $L$ increases. The critical point is obtained by employing a finite-size scaling (FSS) analysis around $g_{AA}=2$, assuming $\overline{\text{IPR}}\propto L^{-D_{f}}$, as presented in Fig. 3 (b), from which one may notice that any $g_{AA}>2$ leads to a finite $\overline{\text{IPR}}$. Similarly, one may check the behavior of $\text{IPR}_{min}$ and its FSS analysis, as shown in Figs. 3 (c), and (d), respectively. One obtains a finite response to $\text{IPR}_{min}$ at $g_{AA}=2$. The former and latter results are consistent with our expectation that in the AAH model all eigenstates become localized at the same critical point. The analysis of the fractal dimension is also consistent with the previous ones, providing $D_{f}=0.53(1)$ at $g_{AA}=2.0$.
Figure 4: Inverse participation ratio results of the AAH model coupled to the electromagnetic field. Here, we fixed $\gamma=0.05$ and $\omega_{0}=1$. (a) $\overline{\text{IPR}}$ as a function of $g_{AA}$ for different system sizes $L$. (b) $\overline{\text{IPR}}$ as a function of $1/L$ (symbols) and extrapolated to the thermodynamic limit (solid lines). (c) IPRmin as a function of $g_{AA}$ for different system sizes $L$. (d) IPRmin as a function of $1/L$ (symbols) and extrapolated to the thermodynamic limit (solid lines).
@@
We now turn to examine the cavity-coupled case. For instance, Fig. 2 (b) displays the energy spectrum and the IPR behavior for fixed $\omega_{0}=1$ and $\gamma=0.15$, from which differences with respect to the no-cavity case are apparent. First, the electron band is broadened due to the contribution of photons, resulting in an almost continuous spectrum. Second, although we can find extended (localized) states for $g_{AA}\ll 2$ ($g_{AA}\gg 2$), the IPR behavior is quite noisy around $g_{AA}=2$. This noise is enhanced or suppressed if the coupling $\gamma$ with the electromagnetic field is increased or reduced, respectively.
@@
Figure 5: The phase diagram of the AAH model coupled to the electromagnetic field of the optical cavity, where we have integrated over the entire energy spectrum. Here, we fixed (a) $\omega_{0}=1$ and (b) $\omega_{0}=2$. The triangles denote the boundaries given by the $\overline{\text{IPR}}$, while the diamonds represent those provided by the $\text{IPR}_{min}$. The solid lines are just guides to the eye.
@@
In order to further quantify these results, we analyze $\overline{\text{IPR}}$ (over the entire spectrum) for fixed $\omega_{0}=1$ and $\gamma=0.05$, as displayed in Fig. 4 (a); the corresponding FSS analysis is shown in Fig. 4(b). The picture that emerges is that despite being a small coupling, it suffices to displace the edge of the localized phase to $g_{AA}\approx 1.8$. This indicates that the critical point shifts due to the coupling with the cavity, in stark contrast to the standard (cavity-free) case. On the other hand, by examining $\text{IPR}_{min}$ and its FSS analysis, presented in Figs. 4 (c) and (d), respectively, we find a ‘critical’ point at $g_{AA}=2$, consistent with the standard AAH model [see, e.g., Figs. 3 (c) and (d)].
@@
One key observation drawn from the previous results is the following: while $\overline{\text{IPR}}$ represents the average of all IPRs, indicating when a fraction of the states ceases to be extended, $\text{IPR}_{min}$ quantifies the presence of at least one extended state. Thus, the former marks the boundary of the phase in which all eigenstates are extended, while the latter marks the boundary of the phase in which all eigenstates are localized. We recall that in the absence of a cavity, $\overline{\text{IPR}}$ and $\text{IPR}_{min}$ lead to the same critical point; see Fig. 3.
Figure 6: Same as Fig. 4, but for fixed $\gamma=0.1$ and $\omega_{0}=2$.
@@
By repeating the same procedure outlined above for different values of $\gamma$ and fixed $\omega_{0}=1$, we obtain the phase diagram shown in Fig. 5 (a). As discussed below, the intermediate region (i.e., the one with finite $\overline{\text{IPR}}$, but vanishing $\text{IPR}_{min}$) exhibits multifractal features. In view of this, we define this region as critical, although there are some subtleties involving the eigenstates which we discuss below. At any rate, the main effect of the electromagnetic field is to spread the critical point into a critical region, which broadens with increasing coupling.
@@
Let us now discuss how the photon frequency changes the shape of the phase diagram. Figure 6 exhibits the behavior of $\overline{\text{IPR}}$ and $\text{IPR}_{min}$ for fixed $\omega_{0}=2$ and $\gamma=0.1$. While $\overline{\text{IPR}}$ is pushed to smaller values of $g_{AA}$ (similarly to the previous case), $\text{IPR}_{min}$ is pushed to larger values of $g_{AA}$, thus significantly expanding the intermediate region. Analyses for other values of $\gamma$ lead to the phase diagram presented in Fig. 5(b), from which we see that the range of the intermediate region is broader than for $\omega_{0}=1$. At this point, it is worth mentioning that if $\hbar\omega_{0}\gg W$, with $W=4t$ being the bare electron bandwidth, the absorption or emission of photons becomes unlikely, thus decreasing the intermediate region.
@@
Figure 7: The energy-resolved phase diagram of the AAH model coupled to the electromagnetic field of the optical cavity. Here, we fixed $\omega_{0}=1$ and $\gamma=0.15$. The triangles denote the boundaries given by the $\overline{\text{IPR}}$, while the diamonds represent those provided by the $\text{IPR}_{min}$. The dashed lines are guides to the eye.
@@
At this point, we recall that in certain cases, such as in gate-induced experiments, in which the voltage at the leads – hence the kinetic energy of the electrons – is well-defined. Therefore, we should have at our disposal the extended-localized behavior within specific energy ranges, rather than averaging over the entire spectrum. In view of this, it is worth analyzing $\overline{\text{IPR}}$ and $\text{IPR}_{min}$ at fixed energy ranges. To this end, by generically fixing $\omega_{0}=1$ and $\gamma=0.15$, while varying the energy, $E$  [^1] Here we average $\overline{\text{IPR}}$ and $\text{IPR}_{min}$ over a spectrum width $|\Delta E|=~{}2$ centered at $E$., we determine the critical points for both quantities, as shown in Fig. 7. The behavior for other values of $\gamma$ is similar, differing only by the size of the intermediate region.
@@
There are several key points we should highlight in relation with Fig. 7. First and foremost, one is still able to identify a critical region, although its width depends significantly on the position of the fixed energy interval, being wider for eigenstates with higher energies. Thus, the boundaries in Fig. 5 should be thought of as lower and upper bounds, respectively for $g_{AA}<2$ and $g_{AA}>2$, across the whole energy spectrum. Second, a perturbative treatment of the coupling with the field shows that transitions between states within subspaces with different number of photons are only significant if the unperturbed states are very close in energy. Even though $\gamma=0.15$ leads to nonperturbative effects, the previous argument provides insights into Fig. 7. Specifically, transitions between states in the zero-photon and one-photon subspaces are unlikely for states at the bottom of the energy spectrum, most of which remain in the zero-photon subspace, regardless of the value of $\gamma$. Consequently, for these eigenstates the effect of the coupling is less pronounced, and they are expected to follow the well-known behavior of the standard AAH model. This explains why the intermediate region shrinks to $g_{AA}\approx 2$ as the energy goes to the bottom of the spectrum in Fig. 7. Such behavior should occur for any value of coupling $\gamma$ and photon frequency $\omega_{0}$.
@@
Now we turn to discuss the nature of the electronic states in this intermediate region. It is reasonable to suppose that the coupling with the electromagnetic field would drive the AAH critical point into a critical region. Curiously, a similar qualitative behavior occurs when dealing with spin-orbit coupling, as discussed in Ref. Zhou et al. (2013). Therefore, in order to examine the critical behavior of the eigenstates in the intermediate region, we now perform the multifractal analysis; see Sec. II.3. Here we deal with PBC, considering $L=F_{m}$ as a number of the Fibonacci sequence.
To achieve this goal we start with the analysis of the usual AAH without the cavity coupling, which will serve as a reference when analyzing the effects of coupling with the electromagnetic field. Figure 8(a) displays the behavior of $\overline{\alpha}_{min}(m)$ as a function of $1/m$, for different system sizes, in the absence of photons  [^2] In the present case, since the phase transition of the AAH model occurs at $g_{AA}=2$ for all eigenstates, we have taken $\Delta E$ as the whole spectrum, i.e. we average over all eigenstates. Their extrapolated values are presented in Fig. 8 (b), where one may notice that, even very close to the critical point (e.g., for $g_{AA}=1.98$ or $2.02$) the extrapolations lead to $\overline{\alpha}_{min}\approx 0$ or 1. Only at $g_{AA}=2$ one obtains an intermediate value of $\overline{\alpha}_{min}$, determining the fractal distribution of the wave function in the medium, and emphasizing that this is the only critical point of the standard AAH model.
Figure 8: The multifractal analysis for both (a)-(b) the standard AAH model and (c)-(d) the case coupled with the cavity. (a) $\overline{\alpha}_{\min}$(m) as a function of $1/m$ for different values of $g_{AA}$, close to the critical point. The solid lines denote their extrapolation to the thermodynamic limit. (b) The extrapolated results of $\overline{\alpha}_{\min}$ as a function of $g_{AA}$. Panels (c) and (d) are the same as panels (a) and (b), respectively, but for fixed $\gamma=0.15$ and $\omega_{0}=1$, while averaging over the range $\Delta E=[0,2]$.
@@
Given this, we repeat the same procedure to the case with the cavity, whose results are presented in Figs. 8(c) and (d), for fixed $\omega_{0}=1$ and $\gamma=0.15$, and for different values of $g_{AA}$. As the intermediate region depends on the energy range, here we have defined $\Delta E=[0,2]$. Notice that $\overline{\alpha}_{min}\approx 1$ when $g_{AA}\lesssim 1.2$, while $\overline{\alpha}_{min}\approx 0$ when $g_{AA}\gtrsim 2.0$, indicating extended and localized states for these regions, respectively. Indeed, the thresholds provided by $\overline{\alpha}_{min}$ are in good agreement with the critical points identified through the $\overline{\text{IPR}}$ and $\text{IPR}_{min}$ analyses, presented in Fig. 7. Within the intermediate region, $1.2\lesssim g_{AA}\lesssim 2.0$, we obtain a continuous range of values for $\overline{\alpha}_{min}$, which strongly suggests the presence of critical states. That is, differently from the standard AAH model, where such a fractal eigenstate occurs only at the critical point $g_{AA}=2$, the coupling with the photons leads to a critical region.
As $\overline{\alpha}_{min}$ is obtained by averaging over a given energy range, it is also appropriate to examine its probability distribution, rather than the mean value. Figure 9 displays $P\big{[}\alpha_{min}\big{]}$ for $g_{AA}=1.1$, 1.7 and 2.3, with fixed $\omega_{0}=1$, $\gamma=0.15$, and $\Delta E=[0,2]$. Notice that $P\big{[}\alpha_{min}\big{]}$ exhibits a clear and sharp peak at high values of $\alpha_{min}$ for $g_{AA}=1.1$, consistent with extended states. For $g_{AA}=2.3$, the distribution also exhibits a peak, but at low values of $\alpha_{min}$, as expected for localized states. By contrast, fixing $g_{AA}=1.7$ within the intermediate region, we see that $P\big{[}\alpha_{min}\big{]}$ broadens considerably. That is, $P\big{[}\alpha_{min}\big{]}$ changes continuously as a function of $g_{AA}$, reshaping from a peak at large values of $\alpha_{min}$ (for large $g_{AA}$), to a peak at small values of $\alpha_{min}$ (for small $g_{AA}$), through a broad distribution in the intermediate region. This strongly suggests a mixture of extended, localized, and critical states within the intermediate region.
@@
Given this, we perform the same analysis for the case with the cavity, fixing $\omega_{0}=1$, $\gamma=0.15$, and $\Delta E=[0,2]$. Figure 10 (d)-(f) shows the behavior of $P[\ln({\rm IPR})]$ for $g_{AA}=1.1$, $1.7$, and $2.3$, respectively. Similarly to the standard AAH model, panels (d) and (f) show the expected behavior for the extended and localized regimes, i.e. with peaks at small and large values of $\ln\big(L\times\text{IPR}\big{missing})$, respectively. However, in the intermediate region (e.g., for $g_{AA}=1.7$) the distribution is broader, as shown in Fig. 10 (e). Interestingly, $P[\ln({\rm IPR})]$ in panel (e) seems to exhibit contributions from size-independent peaks at small $\ln\big(L\times\text{IPR}\big{missing})$, and also from those size-dependent at large values of $\ln\big(L\times\text{IPR}\big{missing})$, while having a continuous distribution in between. This aligns with the behavior of $P\big{[}\alpha_{min}\big{]}$, presented in Fig. 9, thus confirming our claim about the occurrence of a mixture of extended, localized, and critical states at the intermediate region.

### v2

@@ III Results
We begin our analysis by discussing the established properties of the AAH model without a cavity, which will serve as a baseline when discussing the effects of the coupling with photons. Figure 3 (a) shows the spectrum of the model for different disorder strengths. Notice that all eigenstates change their IPR near $g_{AA}=2$, as demanded by the self-dual property of the model. To probe the critical region it is more convenient to examine $\overline{\text{IPR}}$ for different system sizes, as shown in Fig. 4 (a), averaged over the entire spectrum. Notice that for $g_{AA}>2$ $\overline{\text{IPR}}$ is large and weakly dependent on the system size, whereas for $g_{AA}<2$ it is small and decreases as $L$ increases. The critical point is obtained by employing a finite-size scaling (FSS) analysis around $g_{AA}=2$, assuming $\overline{\text{IPR}}\propto L^{-D_{f}}$, as presented in Fig. 4 (b), from which one may notice that any $g_{AA}>2$ leads to a finite $\overline{\text{IPR}}$. The $\text{IPR}_{\rm typ}$ has the same behavior of the $\overline{\text{IPR}}$, as emphasized in Figs. 4 (e) and (f). Similarly, one may check the behavior of $\text{IPR}_{min}$ and its FSS analysis, as shown in Figs. 4 (c), and (d), respectively. One obtains a finite response to $\text{IPR}_{min}$ at $g_{AA}=2$. These findings are consistent with our expectation that in the AAH model all eigenstates become localized at the same critical point. The fractal dimension thus estimated is $D_{f}=0.53(1)$ at $g_{AA}=2.0$, in good agreement with Ref. Evangelou and Pichard (2000). It is important to mention that the fractal dimension $D_{f}$ depends on the energy range of the eigenstates. In fact, $D_{f}$ exhibits a distribution around a mean value that is smaller than 1, which shows that all eigenstates exhibit energy-dependent mutifractal features at the critical point. For instance, near the band center $E\approx 0$, the fractal dimension is $D_{f}\approx 0.82$, also in agreement with Ref. Evangelou and Pichard (2000).
[…]
Figure 5: Same as Fig. 4, but for the AAH model coupled to the electromagnetic field, fixing $\gamma=0.1$ and $\omega_{0}=1$. The results for the IPR are averaged over energies $E\leq 6t$.
@@
We now turn to examine the cavity-coupled case. For instance, Fig. 3 (b) displays the energy spectrum and the IPR behavior for fixed $\omega_{0}=1$ and $\gamma=0.15$, from which differences with respect to the no-cavity case are apparent. First, the electron band is broadened due to the contribution of photons, resulting in an almost continuous spectrum. Second, although we can find extended (localized) states for $g_{AA}\ll 2$ ($g_{AA}\gg 2$), the IPR behavior is quite noisy around $g_{AA}=2$. This noise is enhanced or suppressed if the coupling $\gamma$ with the electromagnetic field is increased or reduced, respectively. At this point, a remark should be made. The broadening of the spectrum arises from the presence of different photon-number subspaces, denoted as $\hat{H}_{n,n}$ in Sec. II.1. Each of these subspaces contains an effective copy of the AAH spectrum (with a small renormalization in the hopping integral), shifted by an energy $n\hbar\omega_{0}$. These subspaces are coupled by the transition terms $\hat{H}_{n,m}$ for $n\neq m$. Therefore, when the photon energy $\hbar\omega_{0}$ is smaller than the bare electronic bandwidth $W\approx 4t$, energy levels from different subspaces can become degenerate. Specifically, high-energy states in $\hat{H}_{n,n}$ may become degenerate with lower-energy states in $\hat{H}_{n+1,n+1}$ or $\hat{H}_{n+2,n+2}$. As these states have distinct energy-dependent physical properties (subsumed by their fractal dimensions), their coupling with the cavity may lead to different transport properties in comparison with the uncoupled states. Indeed, as it will be discussed later, there is a clear dependence on the physical properties with the choice of $\omega_{0}$.
@@
Figure 6: The phase diagram of the AAH model coupled to the electromagnetic field of the optical cavity, where we have integrated over energies $E\leq 6t$. Here, we fixed (a) $\omega_{0}=1$ and (b) $\omega_{0}=2$. The triangles denote the boundaries given by the $\text{IPR}_{\rm typ}$, while the diamonds represent those provided by the $\text{IPR}_{min}$.
@@
In order to further quantify these results, we analyze $\overline{\text{IPR}}$ (averaged over energies $E\leq 6t$) for fixed $\omega_{0}=1$ and $\gamma=0.10$, as displayed in Fig. 5 (a); the corresponding FSS analysis is shown in Fig. 5(b). Similar results for the $\text{IPR}_{\rm typ}$ are presented in Figs. 5 (e) and (f). Since $\text{IPR}_{\rm typ}$ involves taking the logarithm of the IPR before averaging, it highlights the contribution of delocalized states when analyzing an inhomogeneous ensemble. This makes $\text{IPR}_{\rm typ}$ a more reliable indicator for estimating the localization transition; thus, hereafter we use it to determine the critical values. Despite this technical nuance, both the $\overline{\text{IPR}}$ and $\text{IPR}_{\rm typ}$ yield a consistent qualitative picture: the boundary of the localized phase shifts to $g_{AA}\approx 1.6$ for $\gamma=0.10$. That is, the critical point shifts due to the coupling with the cavity, in stark contrast to the standard (cavity-free) case.
@@
On the other hand, by examining $\text{IPR}_{min}$, presented in Fig. 5 (c), we find a small shift of it to $g_{AA}>2$ [see, e.g., Figs. 4 (c)]. Figure 5 (d) displays the FSS analysis of the $\text{IPR}_{min}$, which confirms the persistence of extended states up to approximately $g_{AA}=2.05$. One key conclusion drawn from the previous results is the following: while $\text{IPR}_{\rm typ}$ represents the average of all IPRs, indicating when a fraction of the states ceases to be extended, $\text{IPR}_{min}$ quantifies the presence of at least one extended state. Thus, the former marks the boundary of the phase in which all eigenstates are extended, while the latter marks the boundary of the phase in which all eigenstates are localized. We recall that in the absence of a cavity, $\overline{\text{IPR}}$, $\text{IPR}_{\rm typ}$ and $\text{IPR}_{min}$ lead to the same critical point; see Fig. 4.
Figure 7: Same as Fig. 4, but for the AAH model coupled to the electromagnetic field, fixing $\gamma=0.1$ and $\omega_{0}=2$. The results for the IPR are averaged over energies $E\leq 6t$.
@@
By repeating the same procedure outlined above for different values of $\gamma$ and fixed $\omega_{0}=1$, we obtain the phase diagram shown in Fig. 6 (a). As discussed below, we anticipate that the intermediate region – characterized by a finite $\text{IPR}_{\rm typ}$ but a vanishing $\text{IPR}_{min}$ – contains a mixture of extended, localized, and multifractal states. Due to this coexistence of different types of wavefunction behavior, we refer to this region as critical. However, we note that not all states within this region necessarily exhibit fractal features. At any rate, the main effect of the electromagnetic field is to spread the critical point into a inhomogeneous intermediate region, which broadens with increasing coupling.
@@
Let us now discuss how the photon frequency changes the shape of the phase diagram. Figure 7 exhibits the behavior of $\overline{\text{IPR}}$, $\text{IPR}_{\rm typ}$ and $\text{IPR}_{min}$ for fixed $\omega_{0}=2$ and $\gamma=0.1$. While $\overline{\text{IPR}}$ and $\text{IPR}_{\rm typ}$ are pushed to smaller values of $g_{AA}$ (similarly to the previous case), $\text{IPR}_{min}$ is pushed to larger values of $g_{AA}$, thus expanding the intermediate region. Analyses for other values of $\gamma$ lead to the phase diagram presented in Fig. 6 (b), from which we see that the range of the intermediate region is broader than for $\omega_{0}=1$. We recall that increasing the value of $\omega_{0}$ effectively connects high-energy states within the subspace of a given $n=n_{0}$ photons to low-energy states in subspaces with $n>n_{0}$ photons. In particular, when $\hbar\omega_{0}<W$, where $W=4t$ is the bare electronic bandwidth, energy levels from different photon-number subspaces become degenerate, and non-perturbative corrections become relevant, especially at large values of the coupling parameter $\gamma$. Since certain physical quantities – such as the fractal dimension Evangelou and Pichard (2000) and diffusion coefficient Sutradhar et al. (2019) – depend on the energy of the corresponding undressed states, this coupling between high and low-energy states can lead to significant changes, such as the broadening of the intermediate region. In other words, when degeneracies exist within the photon-number subspaces, a higher-energy photon kick can effectively overcome the system’s tendency to become trapped (for some eigenstates), with the electron only localizing at a larger values of $g_{AA}$. On the other hand, if $\hbar\omega_{0}\gg W$, the absorption or emission of photons becomes unlikely, with corrections being perturbative, thus decreasing the size of the intermediate region.
@@
Figure 8: Energy-resolved phase diagrams of the AAH model coupled to the electromagnetic field of the optical cavity for fixed $\omega_{0}=1$, and (a) $\gamma=0.10$ and (b) $\gamma=0.15$. The square symbols denote the boundaries given by the $\text{IPR}_{\rm typ}$, while the diamonds represent those provided by the $\text{IPR}_{min}$. The hatched area are just guides to the eye.
@@
At this point, we recall that in certain cases the kinetic energy of the electrons is well-defined, particularly in gate-induced experiments where the voltage applied to the leads can be tuned. Therefore, we should have at our disposal the extended-localized behavior within specific energy ranges, rather than averaging over a wider range of the spectrum. In view of this, it is worth analyzing $\text{IPR}_{\rm typ}$ and $\text{IPR}_{min}$ at fixed energy ranges. Here, we average both quantities over a spectrum width $|\Delta E|=~2$, centered at a given $E$. By generically fixing $\omega_{0}=1$, we present the energy-resolved phase diagrams in Fig. 8 for (a) $\gamma=0.10$ and (b) $\gamma=0.15$. The behavior for other values of $\gamma$ is similar, differing only by the size of the intermediate region.
@@
There are several key points we should highlight in relation with Fig. 8. First and foremost, one is still able to identify a critical region, although its width depends significantly on the position of the fixed energy interval, being wider for eigenstates with higher energies. Thus, the boundaries in Fig. 6 should be thought of as lower and upper bounds, respectively for $g_{AA}<2$ and $g_{AA}>2$, across the energy spectrum up to $E=6t$. Second, a perturbative treatment of the coupling with the field shows that transitions between states within subspaces with different number of photons are only significant if the unperturbed states are very close in energy. Even though $\gamma=0.10$ or 0.15 lead to nonperturbative effects, the previous argument provides insights into Fig. 8. Specifically, transitions between states in the zero-photon and one-photon subspaces are unlikely for states at the bottom of the energy spectrum, most of which remain in the zero-photon subspace, regardless of the value of $\gamma$. Consequently, for these eigenstates the effect of the coupling is less pronounced, and they are expected to follow the well-known behavior of the standard AAH model. This explains why the intermediate region shrinks to $g_{AA}\approx 2$ as the energy goes to the bottom of the spectrum in Fig. 8. Such behavior should occur for any value of coupling $\gamma$ and photon frequency $\omega_{0}$.
@@
Now we turn to discuss the nature of the electronic states in this intermediate region. It is reasonable to suppose that the coupling with the electromagnetic field would drive the AAH critical point into a critical region. Curiously, a similar qualitative behavior occurs when dealing with spin-orbit coupling, as discussed in Ref. Zhou et al. (2013). Therefore, in order to examine the critical behavior of the eigenstates in the intermediate region, we now perform the multifractal analysis; see Sec. II.4. Here we deal with PBC, considering $L=F_{m}$ as a number of the Fibonacci sequence.
[…]
@@
Given this, we repeat the same procedure to the case with the cavity, whose results are presented in Figs. 9 (c) and (d), for fixed $\omega_{0}=1$ and $\gamma=0.15$, and for different values of $g_{AA}$. As the intermediate region depends on the energy range, here we have defined $\Delta E=[0,2]$. Notice that $\overline{\alpha}_{min}\approx 1$ when $g_{AA}\lesssim 1.2$, while $\overline{\alpha}_{min}\approx 0$ when $g_{AA}\gtrsim 2.0$, indicating extended and localized states for these regions, respectively. Indeed, the thresholds provided by $\overline{\alpha}_{min}$ are in good agreement with the critical points identified through the $\text{IPR}_{\rm typ}$ and $\text{IPR}_{min}$ analyses, presented in Fig. 8. Within the intermediate region, $1.2\lesssim g_{AA}\lesssim 2.0$, we obtain a continuous range of values for $\overline{\alpha}_{min}$, which strongly suggests the presence of critical states. That is, differently from the standard AAH model, where such a fractal eigenstate occurs only at the critical point $g_{AA}=2$, the coupling with the photons leads to a critical region.
@@
As $\overline{\alpha}_{min}$ is obtained by averaging over a given energy range, it is also appropriate to examine its probability distribution, rather than the mean value. Figure 10 displays $P\big{[}\alpha_{min}\big{]}$ for $g_{AA}=1.1$, 1.7 and 2.3, with fixed $\omega_{0}=1$, $\gamma=0.15$, and $\Delta E=[0,2]$. Notice that $P\big{[}\alpha_{min}\big{]}$ exhibits a clear and sharp peak at large values of $\alpha_{min}$ for $g_{AA}=1.1$, consistent with extended states. For $g_{AA}=2.3$, the distribution also exhibits a peak, but at small values of $\alpha_{min}$, as expected for localized states. By contrast, fixing $g_{AA}=1.7$ within the intermediate region, we see that $P\big{[}\alpha_{min}\big{]}$ broadens considerably. That is, $P\big{[}\alpha_{min}\big{]}$ changes continuously as a function of $g_{AA}$, reshaping from a peak at large values of $\alpha_{min}$ (for large $g_{AA}$), to a peak at small values of $\alpha_{min}$ (for small $g_{AA}$), through a broad distribution in the intermediate region. This strongly suggests a mixture of extended, localized, and critical states within the intermediate region.
@@
Given this, we perform the same analysis for the case with the cavity, fixing $\omega_{0}=1$, $\gamma=0.15$, and $\Delta E=[0,2]$. Figure 11 (d)-(f) shows the behavior of $P[\ln({\rm IPR})]$ for $g_{AA}=1.1$, 1.71.71.7, and 2.32.32.3, respectively. Similarly to the standard AAH model, panels (d) and (f) show the expected behavior for the extended and localized regimes, i.e. with peaks at small and large values of $\ln\big(L\times\text{IPR}\big{missing})$, respectively. However, in the intermediate region (e.g., for $g_{AA}=1.7$) the distribution is broader, as shown in Fig. 11 (e). Interestingly, $P[\ln({\rm IPR})]$ in panel (e) seems to exhibit contributions from size-independent peaks at small $\ln\big(L\times\text{IPR}\big{missing})$, and also from those size-dependent at large values of $\ln\big(L\times\text{IPR}\big{missing})$, while having a continuous distribution in between. This aligns with the behavior of $P\big{[}\alpha_{min}\big{]}$, presented in Fig. 10, thus confirming our claim about the occurrence of a mixture of extended, localized, and critical states in the intermediate region.